the resistance of a platinum wire at o degree celsius is 4 ohm. what will be the resistance at 100 degree celsius, if the temprature coefficient of resistance is 0.0038 degree celsius.
 

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
As we know
 
Rt = Ro (1 + alpha* t)
 
Now
R(100) = R(0) ( 1+ 0.0038 *100)
 
R(100) = 4 * 1.38 = 5.52 ohm